Skip to content
CalcGospel 國際數學圖譜
返回

CIE 9709 2022 November Paper 53 Q4

A Level / CIE / S1

CIE 9709 2022 Nov Paper 53 Paper · Question 4

题目

Problem

Three fair 4-sided spinners each have sides labelled 1, 2, 3, 4. The spinners are spun at the same time and the number on the side on which each spinner lands is recorded. The random variable XX denotes the highest number recorded.

(a) Show that P(X=2)=764\mathrm{P}(X=2)=\frac{7}{64}.

(3)

(b) Complete the probability distribution table for XX.

(2)

xx1234
P(X=x)\mathrm{P}(X=x)764\frac{7}{64}1964\frac{19}{64}

On another occasion, one of the fair 4-sided spinners is spun repeatedly until a 3 is obtained. The random variable YY is the number of spins required to obtain a 3.

(c) Find P(Y=6)\mathrm{P}(Y=6).

(1)

(d) Find P(Y>4)\mathrm{P}(Y>4).

(2)

题目中文翻译

三个均匀的四面陀螺,每个陀螺的各面分别标有 1、2、3、4。同时旋转这三个陀螺,并记录每个陀螺停止时朝下那一面上的数字。随机变量 XX 表示所记录数字中的最大值。

(a) 证明 P(X=2)=764\mathrm{P}(X=2)=\frac{7}{64}。

(3)

(b) 补全 XX 的概率分布表。

(2)

xx1234
P(X=x)\mathrm{P}(X=x)764\frac{7}{64}1964\frac{19}{64}

另一次,将其中一个均匀的四面陀螺重复旋转,直到得到 3 为止。随机变量 YY 表示得到 3 所需的旋转次数。

(c) 求 P(Y=6)\mathrm{P}(Y=6)。

(1)

(d) 求 P(Y>4)\mathrm{P}(Y>4)。

(2)

原卷通用作答要求

Answer all questions. You should use a calculator where appropriate.

You must show all necessary working clearly; no marks will be given for unsupported answers from a calculator.

Give non-exact numerical answers correct to 3 significant figures, or 1 decimal place for angles in degrees, unless a different level of accuracy is specified in the question.

请回答所有题目。适当时应使用计算器。必须清楚展示所有必要的计算过程;仅给出计算器答案而没有过程,不予评分。除非题目另有精度要求,非精确数值答案保留 3 位有效数字,以度为单位的角度保留 1 位小数。

解答

(a)

解法一

思路

展开

先列最大值恰2的七个等概率有序结果,除以全部64种。

答题过程

展开

The maximum is 2 for exactly these ordered spinner outcomes:

(2,2,2), (2,1,1), (1,2,1), (1,1,2),(2,2,2),\ (2,1,1),\ (1,2,1),\ (1,1,2), (2,2,1), (2,1,2), (1,2,2).(2,2,1),\ (2,1,2),\ (1,2,2).

Each has probability 143\frac1{4^3}, so

P(X=2)=743=764.\mathrm{P}(X=2)=\frac7{4^3}=\boxed{\frac7{64}}.

解法二

思路

展开

另一条路线用最大值至多2减去最大值为1。

答题过程

展开

All three scores are at most 2 with probability (24)3\bigl(\frac24\bigr)^3. The maximum is 1 precisely when all three scores are 1. Hence

P(X=2)=(24)3−(14)3=8−164=764.\mathrm{P}(X=2)=\biggl(\frac24\biggr)^3-\biggl(\frac14\biggr)^3 =\frac{8-1}{64}=\boxed{\frac7{64}}.

(b)

解法一

思路

展开

最大值为1必须全1,最大值4用所有点数至多3的补集。

答题过程

展开 P(X=1)=(14)3=164,\mathrm{P}(X=1)=\biggl(\frac14\biggr)^3=\frac1{64}, P(X=4)=1−(34)3=3764.\mathrm{P}(X=4)=1-\biggl(\frac34\biggr)^3=\frac{37}{64}.

Thus

xx11223344
P(X=x)\mathrm{P}(X=x)164\frac1{64}764\frac7{64}1964\frac{19}{64}3764\frac{37}{64}

(c)

解法一

思路

展开

首次3在第6次,前五次非3。

答题过程

展开 P(Y=6)=(34)514=2434096.\mathrm{P}(Y=6)=\biggl(\frac34\biggr)^5\frac14=\boxed{\frac{243}{4096}}.

(d)

解法一

思路

展开

超过4次才首次成功,等价于前4次均非3。

答题过程

展开 P(Y>4)=(34)4=81256.\mathrm{P}(Y>4)=\biggl(\frac34\biggr)^4=\boxed{\frac{81}{256}}.